# Analysis of Shuffling Beyond Pure Local Differential Privacy

### v2

@@ Abstract
Shuffling is a powerful way to amplify privacy of a local randomizer in private distributed data analysis, but existing analyses mostly treat the local differential privacy (DP) parameter $\varepsilon_{0}$ as the only knob and give generic upper bounds that can be loose and do not even characterize how shuffling amplifies privacy for basic mechanisms such as the Gaussian mechanism. We revisit the privacy blanket bound of Balle et al. (the blanket divergence) and develop an asymptotic analysis that applies to a broad class of local randomizers under mild regularity assumptions, without requiring pure local DP. Our key finding is that the leading term of the blanket divergence depends on the local mechanism only through a single scalar parameter $\chi$, which we call the shuffle index. By applying this asymptotic analysis to both upper and lower bounds, we obtain a tight band for $\delta_{n}$ in the shuffled mechanism’s $(\varepsilon_{n},\delta_{n})$-DP guarantee. Moreover, we derive a simple structural necessary and sufficient condition on the local randomizer under which the blanket-divergence-based upper and lower bounds coincide asymptotically. $k$-$R$ families with $k\geq 3$ satisfy this condition, while for generalized Gaussian mechanisms the condition may not hold but the resulting band remains tight. Finally, we complement the asymptotic theory with an FFT-based algorithm for computing the blanket divergence at finite $n$, which offers rigorously controlled relative error and near-linear running time in $n$, providing a practical numerical analysis for shuffle DP.
@@ 1 Introduction
Shuffling has emerged as one of the most powerful tools for improving the utility–privacy trade-off of differentially private (DP) distributed data analysis Liew et al. (2022); Luo et al. (2025). In the single-message model, a population of users first applies a local randomizer to their data and then submits the resulting messages to an anonymizing shuffler. A growing body of work has demonstrated that, even when the local randomizer satisfies only relatively weak privacy guarantees, shuffling can significantly amplify privacy at the level of the aggregated mechanism, leading to substantially improved accuracy compared with pure local DP Feldman et al. (2022); Balle et al. (2019); Erlingsson et al. (2019).
From both a theoretical and a practical perspective, two complementary ingredients are crucial. On the one hand, we need general analytic results that explain how and to what extent shuffling amplifies privacy for a given local mechanism. On the other hand, practitioners ultimately require numerically accurate finite-$n$ estimates of the privacy parameters $(\varepsilon,\delta)$ for concrete protocols. In the central-DP setting, the latter need has led to a shift from coarse composition theorems to tight numerical accountants (for example, for DP-SGD), which in turn has enabled substantially stronger utility–privacy trade-offs in practice Gopi et al. (2021). In contrast, shuffle DP currently lacks, both from a theoretical and a practical perspective, tools that offer a comparable level of accuracy and flexibility.
Limitations of existing shuffle analyses.
Most existing theoretical and numerical analyses of shuffling focus on local randomizers that satisfy pure local DP (i.e., $(\varepsilon_{0},0)$-local DP) and treat $\varepsilon_{0}$ as the main knob controlling privacy amplification (see, e.g., Balle et al. (2019); Erlingsson et al. (2019)). This $\varepsilon_{0}$-centric viewpoint has two important limitations.
@@
First, the parameter $\varepsilon_{0}$ summarizes the local privacy guarantee, but it is a rather crude descriptor from the viewpoint of shuffle DP: it largely ignores the structural properties of the local mechanism that actually govern privacy amplification. In particular, existing general bounds typically do not distinguish between, say, a Laplace mechanism with a given noise scale, a $k$-randomized response ($k$-$R$) mechanism with a given $k$, or other natural families of local mechanisms. As a consequence, generic upper bounds expressed solely in terms of $\varepsilon_{0}$ can be quite loose. Indeed, several works have shown that by exploiting the specific structure of a given mechanism one can obtain significantly tighter bounds on the effect of shuffling (e.g., for $k$-$R$ one can show that privacy amplification becomes stronger as $k$ increases, even when $\varepsilon_{0}$ is kept fixed Feldman et al. (2022); Biswas et al. (2024)). Yet, we currently lack a general understanding of what intrinsic properties of a local mechanism govern its “shuffle efficiency.”
@@
Second, the restriction to pure local DP mechanisms excludes a wide range of natural and practically relevant local randomizers. Pure local DP is a convenient sufficient condition for privacy amplification by shuffling, but it is far from necessary. In practice, one often encounters approximate-DP local mechanisms or mechanisms that do not satisfy any DP condition. Surprisingly, even for the Gaussian mechanism, arguably one of the most prominent mechanisms in central DP, it is not known how to precisely characterize privacy amplification by shuffling. Balle et al. Balle et al. (2019) already pointed out that the Gaussian case is technically challenging, and subsequent work by Koskela et al. Koskela et al. (2022) observed that even numerical accounting for Gaussian shuffling is nontrivial. To date, existing results for shuffling of Gaussian mechanisms are essentially limited to lower bounds Liew and Takahashi (2022); Chua et al. (2024). While there are attempts to generalize shuffling guarantees to approximate local DP (i.e., $(\varepsilon_{0},\delta_{0}>0)$-local DP) in a mechanism-agnostic fashion Feldman et al. (2022); Cheu et al. (2019), such general bounds tend to be overly pessimistic; they often show little visible amplification for moderate values of $n$, because they are driven by worst-case mechanisms that do not reflect the structure of the specific mechanism.
From a practical standpoint, this is problematic: insisting on pure local DP may force system designers to forgo local mechanisms with superior utility properties simply because their analytic and numerical tools cannot analyze them under shuffling.
@@
A central tool in the analysis of shuffle DP is the privacy blanket introduced by Balle et al. Balle et al. (2019). Given a local randomizer $\mathcal{R}$, the privacy blanket is a data-independent distribution $\mathcal{R}_{\mathrm{BG}}$ defined using $\mathcal{R}$. Balle et al. showed that $\delta$ of the shuffled mechanism can be upper-bounded by a divergence computed under the blanket distribution; we refer to this quantity as the blanket divergence. This blanket-divergence bound is currently the best-known general upper bound on privacy amplification by shuffling Su et al. (2025).
However, the blanket divergence itself is still poorly understood, largely due to its analytical and numerical intractability. In particular, Balle et al. Balle et al. (2019) mainly derive generic upper bounds via Hoeffding- and Bennett-type concentration inequalities based on the local DP parameter $\varepsilon_{0}$. Consequently, these bounds can be quite loose and fail to capture how the blanket divergence depends on the structure of the underlying local randomizer, as well as how tight the blanket-based upper bounds are compared with known lower bounds.
This is partly because Balle et al. explicitly aim to obtain finite-$n$ upper bounds. More generally, carrying out a precise finite-$n$ analysis of shuffling through the blanket divergence is intrinsically challenging: DP requires a worst-case guarantee over all neighboring datasets (i.e., taking the supremum of the blanket divergence over all neighboring input pairs $(x_{1},x_{1}^{\prime})$). Identifying the pairs that maximize the blanket divergence at finite $n$ is highly nontrivial in general. In contrast, in the $\varepsilon_{0}$-local DP setting, the single parameter $\varepsilon_{0}$ already summarizes the worst case over all neighboring input pairs, so concentration-based bounds can be expressed directly in terms of $\varepsilon_{0}$ without explicitly resolving the supremum over $(x_{1},x_{1}^{\prime})$.
@@
In light of the above limitations, our goal is to circumvent the apparent finite-$n$ bottlenecks of generic concentration-based analyses by taking a different route, namely an asymptotic analysis of the blanket divergence. While this sacrifices exactness at any fixed $n$, it allows us to address the conceptual and practical limitations discussed above.
On the theoretical side, we replace the $\varepsilon_{0}$-centric view by a single scalar parameter, the shuffle index $\chi$, which encodes the structure of the local randomizer. As shown in Lemma 3.2, by applying the central limit theorem (i.e., normal approximation), for a sequence $\varepsilon_{n}\to 0$ in the regime [^1] This corresponds to the moderate deviation regime required to apply the asymptotic expansion of the central limit theorem. $\varepsilon_{n}=\omega(n^{-1/2})$ and $\varepsilon_{n}=O(\sqrt{\log n/n})$, the blanket divergence for a fixed distinct pair $(x_{1},x_{1}^{\prime})$ admits the asymptotic expansion
@@
where $\varphi$ denotes the PDF of the standard normal distribution and the shuffle index $\chi$ takes the simple form $\sqrt{\gamma}/\sigma$ associated with the pair $(x_{1},x_{1}^{\prime})$. That is, $\chi$ is the only mechanism-dependent constant appearing in the leading term. Thus $\chi$ fully governs how well a given local randomizer interacts with shuffling, in a way that is both quantitatively precise and conceptually transparent.
Moreover, by analyzing the worst-case value of the shuffle index over neighboring input pairs (i.e., the infimum of $\chi(x_{1},x_{1}^{\prime})$ over $x_{1}\neq x_{1}^{\prime}$), we obtain an asymptotic characterization of the worst-case blanket divergence and hence of the shuffled mechanism’s $(\varepsilon_{n},\frac{\alpha}{n}(1+o(1)))$ asymptotic profile (Theorem 3.3) of DP (i.e., upper bound). Because we can express not only upper bounds but also lower bounds in terms of the blanket divergence, we can quantitatively assess the tightness of the upper bound. In particular, we derive a simple structural necessary and sufficient condition on the local randomizer under which the blanket divergence yields an asymptotically optimal characterization (i.e., matches the upper bound and the lower bound) of shuffling (Theorem 3.4). This condition is satisfied, for example, by the $k$-$R$ family with $k\geq 3$. Moreover, we show that the Laplace and Gaussian mechanisms do not satisfy the condition but the upper and lower bounds differ from the true shuffled privacy $\varepsilon$ by at most about $30\%$. This shows that the blanket divergence, when analyzed through the shuffle index, provides a remarkably tight description of shuffling.
On the practical side, asymptotic expansions alone are not sufficient for numerical accounting at finite $n$. We therefore complement our asymptotic analysis with a finite-$n$ algorithm for computing the blanket divergence at the asymptotically worst-case pair $(x_{1},x_{1}^{\prime})$. Our algorithm is based on an FFT approximation of the distribution (Section 4.1), combined with explicit control of the truncation, discretization, and aliasing errors (Theorem 5). Leveraging the asymptotic expansion from Lemma 3.2, we show that the parameters can be tuned by a parameter $\eta$ so that the relative error is $O(\eta)$, while the running time scales as $\widetilde{O}(n/\eta)$ (Theorem 4.3). In this way, the asymptotic theory not only clarifies the intrinsic “shuffle efficiency” of a local randomizer via the shuffle index, but also guides the design and analysis of practical numerical accountants for shuffle DP.
@@
Another line of work seeks to go beyond the $\varepsilon_{0}$-centric view by exploiting additional structure of the local randomizer, typically in the form of a small number of scalar parameters. Feldman et al. Feldman et al. (2023) introduced a stronger clone paradigm and showed that, for mechanisms such as $k$-$R$, one can parameterize the effect of shuffling using $\varepsilon_{0}$ together with two additional scalar quantities, yielding substantially tighter bounds than $\varepsilon_{0}$-only analyses, although even in the $k$-$R$ case exact optimality is not established. Building on this, Wang et al. Wang et al. (2024, 2025) refine the stronger clone-based analysis, using two scalar parameters, and provide sufficient conditions on the local randomizer under which their bounds become optimal. A subtle point is that the proof of Lemma 4.5 in Wang et al. (2024) implicitly assumes the existence of a nontrivial post-processing map (see the errata of Feldman et al. (2023)). Outside specific mechanism classes such as $k$-$R$ where no such map is needed (see Theorem 7.1 of Feldman et al. (2023)), the results should therefore be interpreted with some care.
@@
  • We empirically validate our asymptotic and FFT-based analysis, and in a distribution-estimation task we show that generalized Gaussian mechanisms achieve favorable privacy-accuracy trade-offs compared with pure local DP mechanisms.
@@ 2.1 Differential Privacy (DP)
$\sup_{x_{1:n}\simeq x^{\prime}_{1:n}}\mathcal{D}_{e^{\varepsilon}}\bigl(\mathcal{M}(x_{1:n})\,\big\|\,\mathcal{M}(x^{\prime}_{1:n})\bigr)\leq\delta.$
Sometimes, we say that $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-DP for $x_{1:n}\simeq x^{\prime}_{1:n}$ if $\mathcal{D}_{e^{\varepsilon}}\bigl(\mathcal{M}(x_{1:n})\,\big\|\,\mathcal{M}(x^{\prime}_{1:n})\bigr)\leq\delta$ Sommer et al. (2019).
@@
$\delta(x_{1:n},x_{1:n}^{\prime})$ $\leq\mathcal{D}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{BG}},\gamma}^{\mathrm{blanket}}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}}):=\frac{1}{n\gamma}\mathbb{E}_{\begin{subarray}{c}Y_{1:M}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{BG}}\\ M\sim\mathrm{Bin}(n,\gamma)\end{subarray}}\left[\max\left(\sum_{i=1}^{M}l_{\varepsilon}(Y_{i}),0\right)\right].$

### v3

@@ Abstract
Shuffling is a powerful way to amplify privacy of a local randomizer in private distributed data analysis. Most existing analyses of how shuffling amplifies privacy are based on the pure local differential privacy (DP) parameter $\varepsilon_{0}$. This paper raises the question of whether $\varepsilon_{0}$ adequately captures the privacy amplification. For example, since the Gaussian mechanism does not satisfy pure local DP for any finite $\varepsilon_{0}$, does it follow that shuffling yields weak amplification? To solve this problem, we revisit the privacy blanket bound of Balle et al. (the blanket divergence) and develop a direct asymptotic analysis that bypasses $\varepsilon_{0}$. Our key finding is that, asymptotically, the blanket divergence depends on the local mechanism only through a single scalar parameter $\chi$ and that this dependence is monotonic. Therefore, this parameter serves as a proxy for shuffling efficiency, which we call the shuffle index. By applying this analysis to both upper and lower bounds of the shuffled mechanism’s privacy profile, we obtain a band for its privacy guarantee through shuffle indices. Furthermore, we derive a simple structural, necessary and sufficient condition on the local randomizer under which this band collapses asymptotically. $k$-$R$ families with $k\geq 3$ satisfy this condition, while for generalized Gaussian mechanisms the condition may not hold but the resulting band remains tight. Finally, we complement the asymptotic theory with an FFT-based algorithm for computing the blanket divergence at finite $n$, which offers rigorously controlled relative error and near-linear running time in $n$, providing a practical numerical analysis for shuffle DP.
@@ 1 Introduction
Local privacy Duchi et al. (2013), in which each user applies a local randomizer to their data before sending it, removes the need for a trusted curator and thus enables distributed private data analysis. However, the price of eliminating trust is often a substantial loss in accuracy. Shuffling, viewed as an anonymization layer that breaks the link between users and their messages, offers a compelling middle ground. Since it only anonymizes messages across users, it preserves the utility of the analytics tasks already supported in the local model, such as frequency estimation, histograms/heavy hitters, and mean estimation. Moreover, a growing line of work shows that this simple step can dramatically amplify privacy, allowing protocols to approach trusted-curator privacy-utility trade-offs Feldman et al. (2022); Balle et al. (2019); Erlingsson et al. (2019). Given its practical utility, shuffling has become a key primitive for private distributed data collection and analytics Liew et al. (2022); Luo et al. (2025).
From a practical data-management perspective, we need precise analysis of how and to what extent shuffling amplifies privacy to choose the best local randomizer for a given task. However, existing tools are local randomizer-agnostic and largely tailored to the pure local differential privacy (DP) setting Balle et al. (2019); Erlingsson et al. (2019) (i.e., $(\varepsilon_{0},0)$-local DP) and this $\varepsilon_{0}$-centric viewpoint has two important limitations.
@@
First, the parameter $\varepsilon_{0}$ summarizes the local privacy guarantee, but it is a rather crude descriptor from the viewpoint of shuffling: it largely ignores the structural properties of the local mechanism that actually govern its privacy amplification. In particular, existing general bounds typically do not distinguish between, say, a Laplace mechanism with a given noise scale, a $k$-randomized response ($k$-$R$) mechanism with a given $k$, or other natural families of local mechanisms. As a consequence, generic upper bounds expressed solely in terms of $\varepsilon_{0}$ can be quite loose. Indeed, several works have shown that by exploiting the specific structure of a given mechanism one can obtain significantly tighter bounds on the effect of shuffling (e.g., for $k$-$R$ one can show that privacy amplification becomes stronger as $k$ increases, even when $\varepsilon_{0}$ is kept fixed Feldman et al. (2022); Biswas et al. (2024)). Yet, we currently lack a general understanding of what intrinsic properties of a local mechanism govern its “shuffle efficiency.”
@@
Second, the restriction to pure local DP mechanisms excludes a wide range of natural and practically relevant local randomizers. Pure local DP is a convenient sufficient condition for privacy amplification by shuffling, but it is far from necessary. In practice, one often encounters approximate-DP local mechanisms or mechanisms that do not satisfy any DP condition. Surprisingly, even for the Gaussian mechanism, arguably one of the most prominent mechanisms in central DP, it is not known how to precisely characterize privacy amplification by shuffling. Balle et al. Balle et al. (2019) already pointed out that the Gaussian case is technically challenging, and subsequent work by Koskela et al. Koskela et al. (2022) observed that numerical accounting for Gaussian shuffling is nontrivial. As a result, to date, existing results for shuffling of Gaussian mechanisms are essentially limited to lower bounds Liew and Takahashi (2022); Chua et al. (2024). While there are attempts to generalize shuffling analysis to approximate local DP (i.e., $(\varepsilon_{0},\delta_{0}>0)$-local DP) in a mechanism-agnostic fashion Feldman et al. (2022); Cheu et al. (2019), such general bounds tend to be overly pessimistic; they often show little visible amplification for moderate values of $n$, because they are driven by worst-case mechanisms that do not reflect the structure of the specific mechanism.
@@
A central tool in the analysis of shuffle DP is the privacy blanket introduced by Balle et al. Balle et al. (2019). Given a local randomizer $\mathcal{R}$, the blanket distribution is a data-independent distribution $\mathcal{R}_{\mathrm{BG}}$ defined using $\mathcal{R}$. Balle et al. showed that the privacy profile $\delta(\varepsilon)$ Balle et al. (2018) of the shuffled mechanism can be upper-bounded by a divergence computed under the blanket distribution; we refer to this quantity as the blanket divergence. This blanket-divergence bound is currently the best-known general upper bound on privacy amplification by shuffling Su et al. (2025).
However, the blanket divergence itself is still poorly understood, largely due to its analytical and numerical intractability. In particular, Balle et al. Balle et al. (2019) mainly derive generic upper bounds via Hoeffding- and Bennett-type concentration inequalities based on $\varepsilon_{0}$. Consequently, these bounds can be quite loose and fail to capture how the blanket divergence depends on the structure of the underlying local randomizer.
This is partly because Balle et al. explicitly aim to obtain finite-$n$ upper bounds. More generally, carrying out a precise finite-$n$ analysis of shuffling through the blanket divergence is intrinsically challenging: DP requires a worst-case guarantee over all neighboring datasets (i.e., establishing a uniform upper bound over all neighboring input pairs). Identifying the pairs that maximize the blanket divergence at finite $n$ is highly nontrivial in general. In contrast, in the $\varepsilon_{0}$-local DP setting, the single parameter $\varepsilon_{0}$ already summarizes the worst case over all neighboring input pairs, so concentration-based bounds can be expressed directly in terms of $\varepsilon_{0}$ without explicitly resolving the supremum over all neighboring input pairs.
@@
In light of the above limitations, our goal is to circumvent the apparent finite-$n$ bottlenecks of generic concentration-based analyses by taking a different route, namely an asymptotic but direct analysis of the blanket divergence. While this sacrifices exactness at any fixed $n$, it is tight and allows us to address the conceptual and practical limitations discussed above.
Concretely, the relevant quantity in blanket divergence can be written as a sum of $n$ i.i.d. random variables, so that its behavior is driven by a central limit theorem (CLT)-type phenomenon. We leverage asymptotic expansions of the CLT Petrov (1975) to obtain a sharp leading-order characterization of the blanket divergence. Formally, in the regime [^1] This corresponds to the moderate deviation regime required to apply the asymptotic expansion of the CLT. $\varepsilon_{n}=\omega(n^{-1/2})$ and $\varepsilon_{n}=O(\sqrt{\log n/n})$, the blanket divergence admits the asymptotic expansion
@@
where $\varphi$ denotes the PDF of the standard normal distribution and $\chi$ is a mechanism-dependent constant easily computed from the structure of the local randomizer. Thus $\chi$ fully governs how well a given local randomizer interacts with shuffling. Crucially, this dependence is monotonic: a larger $\chi$ results in a smaller divergence, thereby implying stronger privacy guarantees. This means that $\chi$ serves as a single-number index of privacy after shuffling; hence, we call it the shuffle index, which substitutes for $\varepsilon_{0}$ as a more refined descriptor of the local randomizer’s shuffle efficiency. In other words, we can choose the best local randomizer for shuffling by choosing the one with the largest shuffle index $\chi$.
Moreover, we obtain an asymptotic characterization of the worst-case blanket divergence and hence the shuffled mechanism’s $(\varepsilon\approx\frac{1}{\chi}\sqrt{\frac{\log n}{n}},\delta\approx\tfrac{\alpha}{n})$ guarantee of DP for any $\alpha>0$. Because we can express not only upper bounds but also lower bounds in terms of the blanket divergence, we can quantitatively assess the tightness of the upper bound. In particular, we derive a simple structural necessary and sufficient condition on the local randomizer under which the blanket divergence yields an asymptotically optimal characterization (i.e., matches the upper bound and the lower bound) of shuffling. This condition is satisfied, for example, by the $k$-$R$ family with $k\geq 3$. Moreover, we show that the Laplace and Gaussian mechanisms do not satisfy the condition but the upper and lower bounds narrow in higher privacy regimes.
On the practical side, asymptotic expansions alone are not sufficient for numerical accounting at finite $n$. We therefore complement our asymptotic analysis with a finite-$n$ algorithm for computing the blanket divergence. Our algorithm is based on an FFT approximation of the distribution, combined with explicit control of the truncation, discretization, and aliasing errors. Leveraging the asymptotic expansion, we show that the parameters can be tuned by a parameter $\eta$ so that the relative error is $O(\eta)$, while the running time scales as $\widetilde{O}(n/\eta)$.
@@
A line of work seeks to go beyond the $\varepsilon_{0}$-centric view by exploiting additional structure of the local randomizer, typically in the form of a small number of scalar parameters. Feldman et al. Feldman et al. (2023) introduced a stronger clone paradigm and showed that, for mechanisms such as $k$-$R$, one can parameterize the effect of shuffling using $\varepsilon_{0}$ together with two additional scalar quantities, yielding substantially tighter bounds than $\varepsilon_{0}$-only analyses, although even in the $k$-$R$ case exact optimality is not established. Building on this, Wang et al. Wang et al. (2024, 2025) refine the stronger clone-based analysis, using two scalar parameters, and provide sufficient conditions on the local randomizer under which their bounds become optimal. A subtle point is that the proof of Lemma 4.5 in Wang et al. (2024) implicitly assumes the existence of a nontrivial post-processing map (see the errata of Feldman et al. (2023)). Outside specific mechanism classes such as $k$-$R$ where no such map is needed (see Theorem 7.1 of Feldman et al. (2023)), the results should therefore be interpreted with some care.
@@
  • We empirically validate our asymptotic and FFT-based analysis, and in a distribution-estimation task we show that generalized Gaussian mechanisms achieve favorable privacy-utility trade-offs compared with pure local DP mechanisms.
@@ 2.1 Differential Privacy (DP)
$\delta_{\mathcal{M}}(\varepsilon)=\sup_{x_{1:n}\simeq x^{\prime}_{1:n}}\mathcal{D}_{e^{\varepsilon}}\bigl(\mathcal{M}(x_{1:n})\,\big\|\,\mathcal{M}(x^{\prime}_{1:n})\bigr)\leq\delta.$
The function $\delta_{\mathcal{M}}(\varepsilon)$ is referred to as the privacy profile of $\mathcal{M}$ Balle et al. (2018). Sometimes, we say that $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-DP for $x_{1:n}\simeq x^{\prime}_{1:n}$ if $\mathcal{D}_{e^{\varepsilon}}\bigl(\mathcal{M}(x_{1:n})\,\big\|\,\mathcal{M}(x^{\prime}_{1:n})\bigr)\leq\delta$ Sommer et al. (2019).
@@
$\delta(x_{1:n},x_{1:n}^{\prime})$ $\leq\mathcal{D}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{BG}},\gamma}^{\mathrm{blanket}}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}}):=\frac{1}{n\gamma}\mathbb{E}_{\begin{subarray}{c}Y_{1:M}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{BG}}\\ M\sim\mathrm{Bin}(n,\gamma)\end{subarray}}\left[\left(\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})\right)_{+}\right],$
where $(\cdot)_{+}$ denotes $\max\{0,\cdot\}$.